Find the rule 526;413;477;364;428;315
1
Expert's answer
2012-11-16T02:30:45-0500
516 - 113 = 413
413 + 64 = 477
477 - 113 = 364
364 + 64 = 428
428 - 113 = 315
So, you firstly subtract 113, secondly add 64